# Assouad spectrum of Gatzouras–Lalley carpets

### v2

@@ 2.2 Differentiability and higher-order phase transitions
Piecewise analyticity follows directly from the piecewise formula given in Section 2.1. Moreover, since the distinct parts correspond to distinct analytic curves and any intersection of distinct analytic curves must have a phase transition of some order, we obtain the formula for $H$.
First, to see Item 1, by standard properties of the concave conjugate, the derivative of $\tau^{*}$ at $\alpha$ exists if and only if $\tau$ is strictly concave at all $t$ for which $\alpha\in\partial\tau(t)$. But $\tau$ fails to be strictly concave at $t$ if and only if $t\in I_{n}^{\circ}$ for a homogeneous part $I_{n}$, in which case $\theta=\theta_{n,\min}=\theta_{n,\max}$, as claimed. Since the curves $\psi_{\underline{j}}$ are affine for a homogeneous column $\underline{j}$ and otherwise strictly concave, and moreover there must be at least one column $\underline{i}$ with $s_{\underline{i}}\leq t_{\min}$, there are at most #$\#\eta(\mathcal{I})-1$ points of non-differentiability.
@@
Otherwise, suppose $\theta\in H_{2}$ and $\theta=\theta_{n,\max}<\theta_{n+1,\min}$. Since $(g^{*})^{\prime\prime}=0$ on $(\phi^{-1}(\theta_{n+1,\min}),\phi^{-1}(\theta_{n,\max}))$ and $(g^{*})^{\prime\prime}$ is uniformly bounded away from 00 on $(\phi^{-1}(\theta_{n,\max}),\phi^{-1}(\theta_{n,\min}))$, and since the parameter change $g^{*}(\phi(\theta))/\phi(\theta)$ is smooth, the second derivative does not exist at $\theta_{n,\max}$. The case $\theta_{n,\max}<\theta_{n+1,\min}=\theta$ is analogous. Otherwise, $\theta=\theta_{\max}$, but again $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is constant on $[\theta_{\max},1)$ so the same argument yields non-existence of the second derivative at $\theta_{\max}$.
Combining these two observations yields Item 2 and Item 3, and in particular that every phase transition has either odd order or order $2$.
@@
We consider a Gatzouras–Lalley carpet with three columns each of width $1/3$, the first of which contains only one map (with arbitrary height less than $1/3$). Let $N\in\operatorname{{\mathbb{N}}}$ be large and let $B\subset\operatorname{{\mathbb{R}}}^{N}$ denote the set of all $N$-tuples $(b_{1},\ldots,b_{N})$ with $30<b_{i}<1/$ 3 and $b_{1}+\cdots+b_{N}<1$. Note that to any pair $(\beta,\tilde{\beta})\in B\times B$ there is an attractor $K=K(\beta,\beta^{\prime})$ where the second column has contraction ratios given by $\beta$, and the third given by $\tilde{\beta}$. Fix $\beta=\tilde{\beta}=\beta_{0}$ for some arbitrary initial choice of $\beta_{0}$ whose entries are not all equal, let $t_{\min}$, $t_{\max}$ be the values corresponding to the carpet $K(\beta_{0},\beta_{0})$, and let $t$ satisfy
@@
Moreover, by the implicit function theorem, for typical choices of $(\beta_{0},\beta_{0})$ and for sufficiently large $N$, the set of parameters in $B\times B$ for which at least the first $k$ derivatives match at $t$ is a non-trivial submanifold of $B\times B$ containing the point $(\beta_{0},\beta_{0})$. But given that the first $k$ derivatives match, the parameters for which the first $k+1$ derivatives match is a proper submanifold, so there must exist parameters $(\beta,\tilde{\beta})$ arbitrarily close to $(\beta_{0},\beta_{0})$ such that precisely the first $k$ derivatives of $\psi$ and $\tilde{\psi}$ agree, but the next derivative does not.
Since the parameters $t_{\min}$ and $t_{\max}$ are continuous functions of $(\beta,\tilde{\beta})$, and since the sign of $\psi-\tilde{\psi}$ must change at $t$ since $k$ is odd, it follows that the function $g$ corresponding to the carpet $K(\beta,\tilde{\beta})$ must have a phase transition of order $k$ at $t$. Finally, since $k\geq 3$, $g^{*}$ is differentiable so the derivative of $g^{*}$ is the inverse of $g^{\prime}$. Since the reparameterisation in terms of $\phi$ is smooth, it follows that $\operatorname{dim}^{\theta}_{\mathrm{A}}K(\beta,\tilde{\beta})$ has a phase transition of order $k$ at $\phi^{-1}(g^{\prime}(t))$.
@@ 2.3 Convexity and concavity
    Constant on the interval $[\theta_{\max},1)$.
@@
Cases Item 1 and Item 3 follow directly from the piecewise formula for the Assouad spectrum given in Section 2.1. The remaining case which requires checking is Item 2. If $I_{m}$ is homogeneous, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly convex in a neighbourhood to the left of $\theta_{\max}$ by Item 1.
Otherwise, assume that $I_{m}$ is inhomogeneous with corresponding column $\underline{j}\in\eta(\mathcal{I})$. Writing $\psi=\psi_{\underline{j}}$ and continuing the computation from the proof of Section 2.1 with the same definitions of $F$ and $\varphi$, for all $\theta\in I_{m}^{\circ}$,
@@
For most of the examples we present in this document, the Assouad spectrum is strictly concave on each non-trivial interval $(\theta_{n,\min},\theta_{n,\max})$. However, in Section 2.4.3, we construct an example with a non-trivial inhomogeneous column and a non-trivial open sub-interval of $(\theta_{n,\min},\theta_{n,\max})$ on which $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly convex.
@@ 2.4.1 Homogeneous carpets
Consider the special case of Gatzouras–Lalley carpets whose columns are all homogeneous, i.e. for each $\underline{j}\in\eta(\mathcal{I})$, there is a unique $b_{\underline{j}}$ so that $b_{i}=b_{\underline{j}}$ for all $i\in\eta^{-1}(\underline{j})$. Recall from Lemma 2.2 that $\psi_{\underline{j}}$ is the affine function
@@
and we recall that $s_{\underline{j}}$ is the dimension of column $\underline{j}$. Hence, the function $g(t)$ is piecewise affine, so for any $\theta\in(0,1)$, $\phi(\theta)\in\partial g(t_{n})$ for some $n=0,\ldots,m$. Equivalently, $\theta_{n}\coloneqq\theta_{n,\min}=\theta_{n,\max}$ for all $n=1,\ldots,m$, so case Item 2 of Section 2.1 never occurs.
We can also derive an explicit formula for case Item 3 of Section 2.1. Since $\psi_{\underline{j}}$ are affine for all $\underline{j}\in\eta(\mathcal{I})$, two functions $\psi_{\underline{i}}$ and $\psi_{\underline{j}}$ either have the same slope (in which case one of them does not appear in the formula at all) or intersect at
@@
A plot of the function $g$ and the Assouad spectrum for a system with four columns (three of which are non-trivial) and three phase transitions is given in Fig. 6.
@@ 2.4.2 An example with three columns and six phase transitions
$\theta_{1,\max}$$\theta_{2,\min}$$\theta_{2,\max}$$\theta_{3,\min}$$\theta$$\operatorname{dim}^{\theta}_{\mathrm{A}}K$
@@
A plot of the Assouad spectrum on a restricted domain is given in Fig. 7.
@@ 2.4.3 Inhomogeneous column with corresponding part convex
The Assouad spectrum corresponding to case Item 2 of Section 2.1 is often concave; in particular, this is the case for the other examples given above. However, this is not necessarily the case in general. To give an explicit example, we consider an IFS consisting of two columns. The first column consists of the single map $T_{1,1}(x,y)=(4x/5,y/1000)$. The second column has maps
@@
where $b_{1}=b_{2}=19/100$ and $b_{3}=\cdots=b_{52}=10^{-20}$, and the $t_{j}$ are chosen so that the IFS is a Gatzouras–Lalley carpet.
The convexity of the Assouad spectrum on a non-trivial open sub-interval of $(\theta_{1,\min},\theta_{1,\max})$ can be easily (albeit tediously) verified using the parametric formula Eq. 5. A plot of the Assouad spectrum is given in Fig. 8.
@@ 2.5.1 Bounds on number of phase transitions
We see from Section 2.2 and Section 2.4 that it is much easier to bound the number of phase transitions when all columns are homogeneous than in the general case, and we ask the following question.
@@ 2.5.2 Assouad spectrum of more general self-affine sets
It is natural to ask about the Assouad spectrum of more general self-affine sets. If the condition $b_{i}<a_{i}$ is weakened to $b_{i}\leq a_{i}$ in Eq. 1 for some (but not all) $i$ then we expect that the main results of this paper will still hold. More generally, it would be worthwhile to calculate the formula for carpets of the form considered by Barański [zbl:1116.28008], or for self-affine sponges in higher dimensions.
@@ 2.5.4 Bi-Lipschitz equivalence
However, the Assouad spectrum does not take into account any columns which do not appear in the spectrum partition, whereas such columns may impact bi-Lipschitz equivalence. Here, the lower spectrum might be useful, since it would also take into account the small columns. More generally, classifying Gatzouras–Lalley carpets up to bi-Lipschitz equivalence is a challenging problem (some partial results are given in [arxiv:2111.00752]), and it would be interesting to investigate what more can be said in this direction using the Assouad and lower spectra.
@@ 2.5.5 Intermediate dimensions of Gatzouras–Lalley carpets
Finally, the intermediate dimensions are another family of dimensions which depend on a parameter $\theta\in(0,1)$, and they lie between Hausdorff and box dimension. A precise formula for the intermediate dimensions of Bedford–McMullen carpets is given in [arxiv:2111.05625], with a highly technical proof. It is likely an even more technically challenging problem to calculate a formula for the intermediate dimensions of all Gatzouras–Lalley carpets. However, in light of Section 2.2 Item 1, it would be interesting to know whether the intermediate dimension function is differentiable for Gatzouras–Lalley carpets whose columns are all inhomogeneous. The intermediate dimensions of Bedford–McMullen carpets with distinct Hausdorff and box dimensions have countably many points of non-differentiability.
@@ 3 A variational formula for the Assouad spectrum
For the remainder of this document, we fix a Gatzouras–Lalley IFS $\{T_{i}\}_{i\in\mathcal{I}}$ with attractor $K$. In this section, we establish a variational formula for the Assouad spectrum. After setting up notation in Section 3.1, in LABEL:ss:var we state the variational formula and give a sketch of the proof. The remaining sections are devoted to the proof of the variational formula. We will use the method of types from large deviations theory; an introduction can be found in [dz2010, §2.1.1].
@@ 3.1.1 Symbolic notation and coding
$[\mathtt{i}]=\{\gamma\in\Omega:\mathtt{i}\preccurlyeq\gamma\}.$
@@ 3.1.2 Notation for probability vectors
Since $10<b_{i}<a_{i}<$ 1 for all $i\in\mathcal{I}$, $\Gamma$ takes values in a compact interval

### v4

@@ 2.2 Differentiability and higher-order phase transitions
Piecewise analyticity follows directly from the piecewise formula given in Section˜2.1. Moreover, since the distinct parts correspond to distinct analytic curves and any intersection of distinct analytic curves must have a phase transition of some order, we obtain the formula for $H$.
First, to see Item˜1, by standard properties of the concave conjugate, the derivative of $\tau^{*}$ at $\alpha$ exists if and only if $\tau$ is strictly concave at all $t$ for which $\alpha\in\partial\tau(t)$. But $\tau$ fails to be strictly concave at $t$ if and only if $t\in I_{n}^{\circ}$ for a homogeneous part $I_{n}$, in which case $\theta=\theta_{n,\min}=\theta_{n,\max}$, as claimed. Since the curves $\psi_{\underline{j}}$ are affine for a homogeneous column $\underline{j}$ and otherwise strictly concave, and moreover there must be at least one column $\underline{i}$ with $s_{\underline{i}}\leq t_{\min}$, there are at most $\eta(\mathcal{I})-1$ points of non-differentiability.
@@
Otherwise, suppose $\theta\in H_{2}$ and $\theta=\theta_{n,\max}<\theta_{n+1,\min}$. Since $(g^{*})^{\prime\prime}=0$ on $(\phi^{-1}(\theta_{n+1,\min}),\phi^{-1}(\theta_{n,\max}))$ and $(g^{*})^{\prime\prime}$ is uniformly bounded away from 0 on $(\phi^{-1}(\theta_{n,\max}),\phi^{-1}(\theta_{n,\min}))$, and since the parameter change $g^{*}(\phi(\theta))/\phi(\theta)$ is smooth, the second derivative does not exist at $\theta_{n,\max}$. The case $\theta_{n,\max}<\theta_{n+1,\min}=\theta$ is analogous. Otherwise, $\theta=\theta_{\max}$, but again $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is constant on $\theta_{\max},1)$ so the same argument yields non-existence of the second derivative at $\theta_{\max}$.
Combining these two observations yields Item˜2 and Item˜3, and in particular that every phase transition has either odd order or order $2$.
@@
We consider a Gatzouras–Lalley carpet with three columns each of width $1/3$, the first of which contains only one map (with arbitrary height less than $1/3$). Let $N\in\operatorname{{\mathbb{N}}}$ be large and let $B\subset\operatorname{{\mathbb{R}}}^{N}$ denote the set of all $N$-tuples $(b_{1},\ldots,b_{N})$ with $0<b_{i}<1/3$ and $b_{1}+\cdots+b_{N}<1$. Note that to any pair $(\beta,\tilde{\beta})\in B\times B$ there is an attractor $K=K(\beta,\beta^{\prime})$ where the second column has contraction ratios given by $\beta$, and the third given by $\tilde{\beta}$. Fix $\beta=\tilde{\beta}=\beta_{0}$ for some arbitrary initial choice of $\beta_{0}$ whose entries are not all equal, let $t_{\min}$, $t_{\max}$ be the values corresponding to the carpet $K(\beta_{0},\beta_{0})$, and let $t$ satisfy
@@
[…]
@@ 2.3 Convexity and concavity
    Constant on the interval $\theta_{\max},1)$.
@@
Cases Item˜1 and Item˜3 follow directly from the piecewise formula for the Assouad spectrum given in Section˜2.1. The remaining case which requires checking is Item˜2. If $I_{m}$ is homogeneous, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly convex in a neighbourhood to the left of $\theta_{\max}$ by Item˜1.
Otherwise, assume that $I_{m}$ is inhomogeneous with corresponding column $\underline{j}\in\eta(\mathcal{I})$. Writing $\psi=\psi_{\underline{j}}$ and continuing the computation from the proof of Section˜2.1 with the same definitions of $F$ and $\varphi$, for all $\theta\in I_{m}^{\circ}$,
@@
For most of the examples we present in this document, the Assouad spectrum is strictly concave on each non-trivial interval $(\theta_{n,\min},\theta_{n,\max})$. However, in Section˜2.4.3, we construct an example with a non-trivial inhomogeneous column and a non-trivial open sub-interval of $(\theta_{n,\min},\theta_{n,\max})$ on which $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly convex.
@@ 2.4.1 Homogeneous carpets
Consider the special case of Gatzouras–Lalley carpets whose columns are all homogeneous, i.e. for each $\underline{j}\in\eta(\mathcal{I})$, there is a unique $b_{\underline{j}}$ so that $b_{i}=b_{\underline{j}}$ for all $i\in\eta^{-1}(\underline{j})$. Recall from Lemma˜2.2 that $\psi_{\underline{j}}$ is the affine function
@@
and we recall that $s_{\underline{j}}$ is the dimension of column $\underline{j}$. Hence, the function $g(t)$ is piecewise affine, so for any $\theta\in(0,1)$, $\phi(\theta)\in\partial g(t_{n})$ for some $n=0,\ldots,m$. Equivalently, $\theta_{n}\coloneqq\theta_{n,\min}=\theta_{n,\max}$ for all $n=1,\ldots,m$, so case Item˜2 of Section˜2.1 never occurs.
We can also derive an explicit formula for case Item˜3 of Section˜2.1. Since $\psi_{\underline{j}}$ are affine for all $\underline{j}\in\eta(\mathcal{I})$, two functions $\psi_{\underline{i}}$ and $\psi_{\underline{j}}$ either have the same slope (in which case one of them does not appear in the formula at all) or intersect at
@@
A plot of the function $g$ and the Assouad spectrum for a system with four columns (three of which are non-trivial) and three phase transitions is given in Fig.˜6.
@@ 2.4.2 An example with three columns and six phase transitions
[…]
@@
A plot of the Assouad spectrum on a restricted domain is given in Fig.˜7.
@@ 2.4.3 Inhomogeneous column with corresponding part convex
The Assouad spectrum corresponding to case Item˜2 of Section˜2.1 is often concave; in particular, this is the case for the other examples given above. However, this is not necessarily the case in general. To give an explicit example, we consider an IFS consisting of two columns. The first column consists of the single map $T_{1,1}(x,y)=(4x/5,y/1000)$. The second column has maps
@@
where $b_{1}=b_{2}=19/100$ and $b_{3}=\cdots=b_{52}=10^{-20}$, and the $t_{j}$ are chosen so that the attractor is a Gatzouras–Lalley carpet.
The convexity of the Assouad spectrum on a non-trivial open sub-interval of $(\theta_{1,\min},\theta_{1,\max})$ can be easily (albeit tediously) verified using the parametric formula Eq.˜5. A plot of the Assouad spectrum is given in Fig.˜8.
@@ 2.5 Further work
2.5.1 Natural interpretation of the concave conjugate
The concave conjugate of the Assouad spectrum has also appeared naturally in another setting, namely local smoothing estimates with restricted times [arxiv:2501.12805]. There, the dual of the Assouad spectrum is referred to as the Legendre–Assouad function. It would be worthwhile to determine the relationship between this function and the column pressure $\tau(t)$ from Eq.˜2.
@@ 2.5.2 Bounds on number of phase transitions
We see from Section˜2.2 and Section˜2.4 that it is much easier to bound the number of phase transitions when all columns are homogeneous than in the general case, and we ask the following question.
@@ 2.5.3 Assouad spectrum of more general self-affine sets
It is natural to ask about the Assouad spectrum of more general self-affine sets. If the condition $b_{i}<a_{i}$ is weakened to $b_{i}\leq a_{i}$ in Eq.˜1 for some (but not all) $i$ then we expect that the main results of this paper will still hold. More generally, it would be worthwhile to calculate the formula for carpets of the form considered by Barański [zbl:1116.28008], or for self-affine sponges in higher dimensions.
@@ 2.5.5 Bi-Lipschitz equivalence
However, the Assouad spectrum does not take into account any columns which do not appear in the spectrum partition, whereas such columns may impact bi-Lipschitz equivalence. Here, the lower spectrum might be useful, since it would also take into account the small columns. More generally, classifying Gatzouras–Lalley carpets up to bi-Lipschitz equivalence is a challenging problem (some partial results are given in [doi:10.1088/1402-4896/ad86f2]), and it would be interesting to investigate what more can be said in this direction using the Assouad and lower spectra.
@@ 2.5.6 Intermediate dimensions of Gatzouras–Lalley carpets
Finally, the intermediate dimensions are another family of dimensions which depend on a parameter $\theta\in(0,1)$, and they lie between Hausdorff and box dimension. A precise formula for the intermediate dimensions of Bedford–McMullen carpets is given in [arxiv:2111.05625], with a highly technical proof. It is likely an even more technically challenging problem to calculate a formula for the intermediate dimensions of all Gatzouras–Lalley carpets. However, in light of Section˜2.2 Item˜1, it would be interesting to know whether the intermediate dimension function is differentiable for Gatzouras–Lalley carpets whose columns are all inhomogeneous. The intermediate dimensions of Bedford–McMullen carpets with distinct Hausdorff and box dimensions have countably many points of non-differentiability.
@@ 3 A variational formula for the Assouad spectrum
For the remainder of this document, we fix a Gatzouras–Lalley IFS $\{T_{i}\}_{i\in\mathcal{I}}$ with attractor $K$. In this section, we establish a variational formula for the Assouad spectrum. After setting up notation in Section˜3.1, in Section˜3.2 we state the variational formula and give a sketch of the proof. The remaining sections are devoted to the proof of the variational formula. We will use the method of types from large deviations theory; an introduction can be found in [zbl:1177.60035, §2.1.1].
@@ 3.1.1 Symbolic notation and coding
$\mathtt{i}]=\{\gamma\in\Omega:\mathtt{i}\preccurlyeq\gamma\}.$
@@ 3.1.2 Notation for probability vectors
Since $0<b_{i}<a_{i}<1$ for all $i\in\mathcal{I}$, $\Gamma$ takes values in a compact interval